OpenStudy (anonymous):

lim as x approaches positive infinity of (1 + 1/4x) ^ 3x

OpenStudy (freckles):

I would use the fact that e^(ln(y))=y

OpenStudy (freckles):

\[\lim_{x \rightarrow \infty}e^{3x \ln(1+\frac{1}{4x})}=\lim_{x \rightarrow \infty}e^{3 \frac{\ln(1+\frac{1}{4x})}{\frac{1}{x}}}\]

OpenStudy (freckles):

now use l'hospitl rule since you have 0/0 above in the exponent

OpenStudy (anonymous):

so the answer would be e^3

OpenStudy (freckles):

I don't think that is correct....

OpenStudy (freckles):

what is the derivative of ln(1+1/(4x)) and what is the derivative of 1/x?

OpenStudy (anonymous):

(1/1+1/4x) * -1/x^2 and -1/x^2

OpenStudy (freckles):

see that is what is going on you are forgetting the constant multiple inside

OpenStudy (freckles):

\[\frac{d}{dx}\ln(1+\frac{1}{4x})=\frac{-1}{4x^2} \frac{1}{1+\frac{1}{4x}} \\ \frac{d}{dx}\frac{1}{x}=\frac{-1}{x^2}\]

OpenStudy (freckles):

\[\lim_{x \rightarrow \infty}e^{3x \ln(1+\frac{1}{4x})}=\lim_{x \rightarrow \infty}e^{3 \frac{\ln(1+\frac{1}{4x})}{\frac{1}{x}}} =\lim_{x \rightarrow \infty}e^{3 \frac{-1}{4x^2} \frac{1}{1+\frac{1}{4x}} \cdot (-x^2)}\]

OpenStudy (anonymous):

so it would result in e^3/4

OpenStudy (freckles):

yes

OpenStudy (anonymous):

awesome!! thank you for your help!!

OpenStudy (freckles):

the x^2's will cancel \[\lim_{x \rightarrow \infty}e^{3x \ln(1+\frac{1}{4x})}=\lim_{x \rightarrow \infty}e^{3 \frac{\ln(1+\frac{1}{4x})}{\frac{1}{x}}} =\lim_{x \rightarrow \infty}e^{3 \frac{1}{4\cancel{-x^2}} \frac{1}{1+\frac{1}{4x}} \cdot (\cancel{-x^2})} \\ =\lim_{x \rightarrow \infty}e^{\frac{3}{4} \frac{1}{1+\frac{1}{4x}}}\] and then that one part goes to 1/(1+0)=1 and then you have 3/4*1=3/4 and blah blah

OpenStudy (freckles):

or in the words of Seinfeld yada yada yada

ganeshie8 (ganeshie8):

or if you're allowerd, you could also use the standard limit \(\large\lim\limits_{n\to\infty} (1+x/n)^n = e^x \)

ganeshie8 (ganeshie8):

freckles work is a nice derivation of above standard result i think..